 1 \ifx\LdfInit\@undefined
 2 \def\LdfInit{%
 3  \chardef\atcatcode=\catcode`\@
 4  \catcode`\@=11\relax
 5  \input babel.def\relax
 6  \catcode`\@=\atcatcode \let\atcatcode\relax
 7  \LdfInit}
 8 \fi
 9 ⟨+american⟩\def\CurrentOption{american}
10 ⟨+british⟩\def\CurrentOption{british}
11 ⟨+english⟩\def\CurrentOption{english}

12 ⟨american | british | english⟩\input english.ldf\relax

13 ⟨+bahasa⟩\input bahasa.ldf\relax
14 ⟨+breton⟩\input breton.ldf\relax
15 ⟨+catalan⟩\input catalan.ldf\relax
16 ⟨+croatian⟩\input croatian.ldf\relax
17 ⟨+czech⟩\input czech.ldf\relax
18 ⟨+danish⟩\input danish.ldf\relax
19 ⟨+afrikaans⟩\def\CurrentOption{afrikaans}
20 ⟨+dutch⟩\def\CurrentOption{dutch}
21 ⟨+afrikaans | dutch⟩\input dutch.ldf\relax
22 ⟨+esperant⟩\input esperant.ldf\relax
23 ⟨+estonian⟩\input estonian.ldf\relax
24 ⟨+finnish⟩\input finnish.ldf\relax
25 ⟨+francais⟩\def\CurrentOption{francais}
26 ⟨+french⟩\def\CurrentOption{french}
27 ⟨+francais | french⟩\input frenchb.ldf\relax
28 ⟨+galician⟩\input galician.ldf\relax
29 ⟨+austrian⟩\def\CurrentOption{austrian}
30 ⟨+german⟩\def\CurrentOption{german}
31 ⟨+germanb⟩\def\CurrentOption{german}

32 ⟨+austrian | german | germanb⟩\input germanb.ldf\relax

33 ⟨+greek⟩\input greek.ldf\relax
34 ⟨+irish⟩\input irish.ldf\relax
35 ⟨+italian⟩\input italian.ldf\relax
36 ⟨+lsorbian⟩\input lsorbian.ldf\relax
37 ⟨+magyar⟩\def\CurrentOption{magyar}
38 ⟨+hungarian⟩\def\CurrentOption{hungarian}
39 ⟨+magyar | hungarian⟩\input magyar.ldf\relax
40 ⟨+norsk⟩\def\CurrentOption{norsk}
41 ⟨+nynorsk⟩\def\CurrentOption{nynorsk}
42 ⟨+norsk | nynorsk⟩\input norsk.ldf\relax
43 ⟨+polish⟩\input polish.ldf\relax
44 ⟨+portuges⟩\def\CurrentOption{portuges}
45 ⟨+portuguese⟩\def\CurrentOption{portuguese}
46 ⟨+brazil⟩\def\CurrentOption{brazil}
47 ⟨+brazilian⟩\def\CurrentOption{brazilian}

48 ⟨+portuges | portuguese | brazil | brazilian⟩\input portuges.ldf\relax

49 ⟨+romanian⟩\input romanian.ldf\relax

50 ⟨+russian⟩\input russianb.ldf\relax
51 ⟨+scottish⟩\input scottish.ldf\relax
52 ⟨+slovak⟩\input slovak.ldf\relax
53 ⟨+slovene⟩\input slovene.ldf\relax
54 ⟨+spanish⟩\input spanish.ldf\relax
55 ⟨+swedish⟩\input swedish.ldf\relax
56 ⟨+turkish⟩\input turkish.ldf\relax
57 ⟨+usorbian⟩\input usorbian.ldf\relax
58 ⟨+welsh⟩\input welsh.ldf\relax